# Give me a hint: Can LLMs take a hint to solve math problems?

### v1

@@ Abstract
While many state-of-the-art LLMs have shown poor logical and basic mathematical reasoning, recent works try to improve their problem-solving abilities using prompting techniques. We propose giving "hints" to improve the language model’s performance on advanced mathematical problems, taking inspiration from how humans approach math pedagogically. We also test the model’s adversarial robustness to wrong hints. We demonstrate the effectiveness of our approach by evaluating various LLMs, presenting them with a diverse set of problems of different difficulties and topics from the MATH dataset and comparing against techniques such as one-shot, few-shot, and chain of thought prompting.
@@ 1 Introduction
The ability to reason and logically solve complex mathematical problems is essential for progress in nearly every field, whether it be modeling complex environments, developing new algorithms, or engineering new devices. Recent works have explored the mathematical capabilities of Large Language models and have found them lacking in logic and basic mathematical reasoning [1] [2]. In this work, we propose "hinting" LLMs, by giving subtle guidance or clues, as a method to improve problem-solving capabilities on mathematical tasks and then compare its effectiveness against other prompting methods, such as one-shot [3], few-shot [3], and chain of thought prompting [4]. We evaluate our approach using the MATH dataset [5], consisting of math problems categorized into distinct classes based on subtopics such as algebra, probability, geometry, etc., with different levels of complexity (1-5). We test a diverse set of LLMs, including base language models, instruction fine-tuned models, models specifically tuned for mathematical tasks, and closed source models such as GPT-4o-mini [6] and Gemini Flash [7].
We further examine robustness to adversarial prompts, misleading hints and clues of varying levels. We investigate how sensitive the models are to incorporating these incorrect hints as context, which may degrade performance, versus their ability to reject such misleading information. Through these experiments, we seek to contribute to the ongoing research on improving current state-of-the-art language model’s reasoning capabilities and their practical applications in solving mathematical tasks[8].
@@ 2 Method
We extracted hints by providing the problems in the MATH dataset to Gemini-1.5-flash model and prompting it to generate a hint for that question. These hints are then provided in context to the model along with the target problem to help the model reason about the task. We believe this is more aligned with how humans solve math problems by getting hints instead of the complete solution for an example problem, as in Chain of Thought, or just the final answer, as in one/few-shot approaches. We also test the adversarial robustness of models to being provided either a adversarial misleading hint, or a hint from a random question. We also generate similar versions for one/few shot and CoT prompting as shown in Figure 1. More details about the prompts can be found in Appendix A.
We evaluate on a diverse ensemble of LLMs ranging from base models to instruct fine-tuned models and math-finetuned LLMs, including nine open-source models and two closed-source models. (More details in C). We use the MATH dataset [5], which contains problems of seven different classes (e.g., algebra, geometry, etc.) of varying difficulty levels from 1 (easy) to 5 (hard). More details about the difficulty levels are given in Appendix B.
@@ 3.1 Setup
We evaluate 11 models for our experiments, using open-source implementations where possible and public API offerings for closed-source models. We evaluate 17 different types of prompting techniques, ranging from baseline to hinting to adversarial, along with one/few shot and CoT as given in Appendix A. The problem set consisted of all seven topics from the MATH dataset [5], with 100 problems for each topic. Exact details about the data split are given in Appendix B. We report the comparison by checking the fraction of question that the model got correct.
@@ 3.2 Evaluating Hint based prompting
We observe that hinting improves the performance of models, as shown in Figure 2. We attribute the poorer performance of other approaches to a lack of generalization. Approaches like chain of thought [4] give step-by-step solutions, which might not generalize to other problems restricting the search space of these models. This can also be due to a "snowball" effect [5], in which intermediate generated steps with mistakes can derail the model from the logical direction to the right answer. One-shot and few-shot perform relatively better as they do not restrict the steps, however, the model still gets confined to a narrower subdomain, with few-shot showing more generalisation due to more examples and hence, better performance. Hinting leads to better performance as it helps the model reason about question-specific knowledge more than other prompting techniques and gives initial steps to ensure a right direction without overfitting to an exact solution.
@@ 3.3 Evaluating Adversarial Hinting
We find that giving Adversarial hints drastically reduces the model’s performance dropping it below CoT which performed the worst in our non-adversarial approaches as shown in Fig 2 and Table 1. We also apply this to few-shot[3] and one-shot[3] settings and find that adversarial hinting affects the performance in those cases as well. Inferior performance of CoT and it’s prompting variants can be attributed to over-fitting to various factors within irrelevant information that influence the model’s sensitivity to distractions based on lexical similarity. [14].
Figure 3: A comparison of various prompting techniques
@@ 3.4 Model-wise performance
While comparing the performance of different models, we observe that closed-source models exhibit the highest performance, followed by models fine-tuned for mathematical tasks, instruct-tuned models and finally base models. Additionally, variations in model size impact performance, with smaller models generally performing worse. However, performance also depends on the extent of fine-tuning; for example, the Qwen-2-Math-Instruct model[15] achieves comparable results to GPT-4o-mini[6] as shown in Appendix E. Among the 7B and 8B models, Qwen-2-Math-Instruct[15] performs best, while Mistral-Instruct[16] ranks the lowest. Our observations further reveal that base models struggle to incorporate and utilize hints effectively, performing worse than the baseline, likely due to their limited ability to follow instructions[17]. We list our exhaustive results in Appendix E.
Further, we also evaluated hinting multimodal models on visual and mathematical tasks and found a similar improvement due to hinting and deterioration due to adversarial hinting. We provide more details in Appendix Section C.1.
@@ 4 Conclusion
In this work, we have evaluated the mathematical reasoning abilities of various models and approaches. Our results indicate that providing hints is more effective than giving direct answers because it guides the model to the correct solution instead of restricting its search space like few-shot[3], one-shot[3], and chain of thought[4] which lack generalization and only perform well when the example and target problems are very similar. Hence, for math problems with unknown solutions, it is better to provide the user’s knowledge about the problem as hints to improve its reasoning and problem-solving capabilities than a full solution for a similar problem. Sensitivity to random and adversarial prompting[18] techniques demonstrated performance loss due to adverserial/random hints in few-shot[3], one-shot[3], and chain-of-thought[4] settings.
@@ 5 Limitations and Future Work
Our work mainly focuses on prompting LLMs with problems as textual input. Testing these techniques in multi-modal models like VLMs, where geometrical problems can be passed as image inputs, is a possible future direction. Further, our experiments involved a small subset of problems to test the models on due to computational limitations.
@@
    Our techniques are yet to be tested on larger models like the Llama 3.1 70B, 450B[19], Falcon 180B[20], OPT 175B[21], etc., and other closed source models like Claude[22], Bard[23], etc due to computational limitations.
@@ Appendix B Dataset
We evaluate on a subset of the MATH dataset [5]. The MATH dataset consists of problems from various popular math competitions, including the AMC 12, AIME, and more. Owing to the structure of mathematics, these problems have a particular method of solving them with multiple related steps. Humans generally use problem-solving techniques and “heuristics” to solve such problems, thus making them a good metric for assessing a model’s problem-solving and reasoning skills.[5] The dataset categorizes the problems into various categories and difficulties. The seven categories are Pre-algebra, Algebra, Number Theory, Counting and Probability, Geometry, Intermediate Algebra, and Pre-calculus. The problems are divided into five levels: level 1, including more straightforward questions; and level 5, including advanced questions.
@@ Appendix C Experiments
All the prompting techniques were tested on a set of models, chosen in manner to ensure diversity for our experimentation’s and testing. Nine open-sourced models and two closed-source have been used models as listed below:
@@ C.1 Experiments on VLMs
We evaluated the mathematical visual question answering ability of multi-modal models by providing them with the text for a question and an image containing the diagram. We tested the visual problem-solving with four prompting techniques: baseline, hinting, adversarial hinting, and random hinting. We evaluated these approaches on the multimodal Gemini and GPT-4o mini models. For our multimodal experiments, we created another dataset using a subset of 700 non-MCQ questions from the Math Vision Dataset[26], which contains problems of various levels from 1-5, in increasing difficulty. We sampled equally from each difficulty level. Hints were generated using the Gemini-1.5-Flash model using the textual content of the question. We also generated their adversarial hints in a similar manner as our other evaluations. For random hinting, we provided the hint from another question. As seen in Table 3, hinting improves the overall model performance, and the performance is degraded due to adversarial and random hinting, similar to our other evaluations. We note that we could not exhaustively evaluate a broader set of multimodal models on a larger dataset due to computational limitations, and see it as an interesting future direction to explore.

### v2

@@ Abstract
While state-of-the-art LLMs have shown poor logical and basic mathematical reasoning, recent works try to improve their problem-solving abilities using prompting techniques. We propose giving "hints" to improve the language model’s performance on advanced mathematical problems, taking inspiration from how humans approach math pedagogically. We also test robustness to adversarial hints and demonstrate their sensitivity to them. We demonstrate the effectiveness of our approach by evaluating various diverse LLMs, presenting them with a broad set of problems of different difficulties and topics from the MATH dataset and comparing against techniques such as one-shot, few-shot, and chain of thought prompting. Our code is available at https://github.com/vlgiitr/LLM-Math
@@ 1 Introduction
The ability to reason and logically solve complex mathematical problems is essential for progress in nearly every field, whether it be modeling complex environments, developing new algorithms, or engineering new devices. Recent works have explored Large Language models’ mathematical capabilities and found them lacking in logic and basic mathematical reasoning [1] [2]. Improving these reasoning capabilities is important in a future where LLMs can be used to assist humans in increasingly complex mathematical tasks or act like AI math teachers. In this work, taking inspiration from how humans are taught math pedagogically, we propose "hinting" LLMs by giving subtle guidance or clues as a method to improve problem-solving capabilities on mathematical tasks and then compare its effectiveness against other prompting methods, such as one-shot [3], few-shot [3], and chain of thought prompting [4]. We evaluate our approach using the MATH dataset [5], consisting of math problems categorized into distinct classes based on subtopics such as algebra, probability, geometry, etc., with different levels of complexity (1-5). We use a diverse set of LLMs, including base language models, instruction fine-tuned models, models specifically tuned for mathematical tasks, and closed source models such as GPT-4o-mini [6] and Gemini Flash [7] for our evaluation.
We further examine the robustness of these models to adversarial prompts, misleading hints, and clues of varying levels. We investigate how sensitive the models are to incorporating these incorrect hints as context, which may degrade performance, versus their ability to reject such misleading information. Through these experiments, we seek to contribute to the ongoing research on improving the current state-of-the-art language model’s reasoning capabilities and their practical applications in solving mathematical tasks [8].
@@ 2 Method
We first prompt the Gemini 1.5 flash model to generate hints for our dataset by passing it the question and final answer. These hints are then provided in context to the model and the target problem to help the model reason about the task. We believe this aligns more with how humans solve math problems by getting hints instead of the complete solution for an example problem, as in Chain of Thought, or just the final answer, as in one/few-shot approaches.
To test the adversarial robustness of these models to hints, we provided either an adversarial misleading hint or a hint from a random question to observe how sensitive the models are to our hints. Similar versions for one/few shot and CoT prompting are also generated as shown in Figure 1. More details about the prompting process can be found in Appendix A.
We evaluate a diverse ensemble of LLMs ranging from base models to instruct fine-tuned models and math-finetuned LLMs, including nine open-source models and two closed-source models (More details in Appendix C). We use the MATH dataset [5], which contains problems of seven different classes (e.g., algebra, geometry, etc.) of varying difficulty levels from 1 (easy) to 5 (hard), more details about the difficulty levels are given in Appendix B.
@@ 3.1 Setup
We evaluate a set of 11 models for our experiments, using open-source implementations where possible and public API offerings for closed-source models. We evaluate 17 different types of prompting techniques, ranging from baseline to hinting to adversarial, along with one/few shot and CoT as given in Appendix A. The problem set consisted of all seven topics from the MATH dataset [5], with 100 problems for each topic. Exact details about the data split are given in Appendix B. We report the comparison by checking the fraction of questions that the model got correct.
@@ 3.2 Evaluating Hint based prompting
We observe that hinting improves the performance of models, as shown in Figure 2. We attribute the poorer performance of other approaches to a lack of generalization. The improvement of hinting over chain of thought, can be explained as the latter restricts the model’s generalization by forcing it to follow the entire reasoning steps of a solution, which might not be the same for all the problems. On the other hand, hints only provide certain helpful directions, giving the model more freedom to generalize better to different problems.
Approaches like the chain of thought [4] give step-by-step solutions, which might not generalize to other problems restricting the search space of these models. This can also be due to a "snowball" effect [5], in which intermediate-generated steps with mistakes can derail the model from the logical direction to the right answer. One-shot and few-shot perform relatively better as they do not restrict the steps, however, the model still gets confined to a narrower subdomain, with few-shot showing more generalisation due to more examples and hence, better performance. Hinting leads to better performance as it helps the model reason about question-specific knowledge more than other prompting techniques and gives initial steps to ensure the right direction without over-fitting to an exact solution.
@@ 3.3 Evaluating Adversarial Hinting
We find that giving Adversarial hints drastically reduces the model’s performance, dropping it below CoT, which performed the worst in our non-adversarial approaches, as shown in Fig 2 and Table 1. We also apply this to few-shot[3] and one-shot[3] settings and find that adversarial hinting affects the performance in those cases as well. The inferior performance of CoT and its prompting variants can be attributed to over-fitting various factors within irrelevant information that influence the model’s sensitivity to distractions based on lexical similarity. [14].
Figure 3: Adversarial and random hinting strategies
@@ 3.4 Model-wise performance
While comparing the performance of different models, we observe that closed-source models exhibit the highest performance, followed by models fine-tuned for mathematical tasks, instruct-tuned models, and finally, base models, as shown in table 2. Additionally, variations in model size impact performance, with smaller models generally performing worse. However, performance also depends on the extent of fine-tuning; for example, the Qwen-2-Math-Instruct model[15] achieves comparable results to GPT-4o-mini[6] as shown in Appendix E. Among the 7B and 8B models, Qwen-2-Math-Instruct[15] performs best, while Mistral-Instruct[16] ranks the lowest. Our observations further reveal that base models struggle to incorporate and utilize hints effectively, performing worse than the baseline, likely due to their limited ability to follow instructions[17]. We list our exhaustive results in Appendix D.
Further, we also evaluated hinting multimodal models on visual and mathematical tasks and found a similar improvement due to hinting and deterioration due to adversarial hinting. We provide more details in Appendix C.1.
@@ 4 Conclusion
In this work, we have evaluated the mathematical reasoning abilities of various models and approaches. Our results indicate that providing hints is more effective than giving direct answers because it guides the model to the correct solution instead of restricting its search space like few-shot[3], one-shot[3], and chain of thought[4] which lack generalization. However, CoT may outperform hinting in some cases where the example and target problem are very similar. Hence, for math problems with unknown solutions, it is better to provide the user’s knowledge about the problem as hints to improve its reasoning and problem-solving capabilities than a full solution for a similar problem, and hinting is the natural way in which humans solve a reasoning task as well, as only intermediate directions to approach a problem are required, the rest can be reasoned by humans themselves. Although extracting hints requires and extra inference, it is compensated by the increase in performance over other prompting techniques, which are only useful when dealing with similar problems. Finally, we see the sensitivity to random and adversarial prompting[18] techniques demonstrated performance loss due to adversarial/random hints in few-shot[3], one-shot[3], and chain-of-thought[4] settings.
@@ 5 Limitations and Future Work
Our work mainly focuses on prompting LLMs with problems as textual input. Testing these techniques in multi-modal models like VLMs, where geometrical problems can be passed as image inputs, is a possible future direction. Further, due to computational limitations, our experiments involved a small subset of problems to test the models. These techniques can also be evaluated on larger sample sizes to increase generalizability. Additionally, It is yet to be tested whether hinting can increase performance in other general tasks as well.
@@
    Our techniques are yet to be tested on larger models like the Llama 3.1 70B, 450B[19], Falcon 180B[20], OPT 175B[21], etc., and other closed source models like Claude Sonnet[22], Bard[23], etc due to computational limitations.
@@ Appendix B Dataset
We evaluate on a subset of the MATH dataset [5]. The MATH dataset consists of problems from various popular math competitions, including the AMC 12, AIME, and more. Owing to the structure of mathematics, these problems have a particular method of solving them with multiple related steps. Humans generally use problem-solving techniques and “heuristics” to solve such problems, thus making them a good metric for assessing a model’s problem-solving and reasoning skills[5]. The dataset categorizes the problems into various categories and difficulties. The seven categories are Pre-algebra, Algebra, Number Theory, Counting and Probability, Geometry, Intermediate Algebra, and Pre-calculus. The problems are divided into five levels: level 1, including more straightforward questions; and level 5, including advanced questions.
@@ Appendix C Experiments
All the prompting techniques were tested on a set of models, chosen in a manner to ensure diversity for our experimentation and testing. Nine open-sourced models and two closed-source have been used models as listed below:
@@ C.1 Experiments on VLMs
We evaluated the mathematical visual question answering ability of multi-modal models by providing them with the text for a question and an image containing the diagram. We tested the visual problem-solving with four prompting techniques: baseline, hinting, adversarial hinting, and random hinting. We evaluated these approaches on the multimodal Gemini and GPT-4o mini models. For our multimodal experiments, we created another dataset using a subset of 700 non-MCQ questions from the Math Vision Dataset[26], which contains problems of various levels from 1-5, in increasing difficulty. We sampled equally from each difficulty level. Hints were generated using the Gemini-1.5-Flash model using the textual content of the question. We also generated their adversarial hints in a similar manner as our other evaluations. For random hinting, we provided the hint from another question. As seen in Table 3, hinting improves the overall model performance, and the performance is degraded due to adversarial and random hinting, similar to our other evaluations. We note that we could not exhaustively evaluate a broader set of multimodal models on a larger dataset due to computational limitations and see it as an interesting future direction to explore.